OpenStudy (anonymous):

You are going to fence an area that is in the shape of a right triangle. The two legs measure 7.8 m and 5.78 m. Estimate the meters of fencing you will need for the longest side, using whole meters and square root.(

OpenStudy (anonymous):

find the third side using pythagorean's theorum. When estimating, use the rules of rounding, in this case, round 7.8 to 8 and 5.78 to 6. the length of the hypotenuse is \[\sqrt{8^{2}+6^{2}}\] or \[\sqrt{64+36}\] = 10 add all three sides together to get an estimated need for 10 + 6 + 8 feet = 24 ft.